# Schrödinger Symmetry in Spherically-symmetric Static Mini-superspaces with Matter Fields

### v1

@@ V.1 Y-coordinate in the mini-superspace
where (diagonal of $\eta^{ab})\equiv(-1,1,\delta_{IJ})$ is a $(2+n)\times(2+n)$ “Lorentzian” constant matrix. This contains only one coordinate $Y^{0}$ so that we can implement the canonical-transformation program easily.
@@ V.2 $(2+n)$ D-Schrödinger symmetry
$\mathcal{G}_{\phi ab}=(-\epsilon)cl_{p}\eta_{ab},~~\mathcal{G}_{\phi}^{ab}=\frac{1}{(-\epsilon)cl_{p}}\eta^{ab},$ (92)
where $\eta_{ab}=\mathrm{diag}(-1,1,\cdots.,1)$. Here, for convenience, we multiplied 1-1 in the right hand side so that the Hamiltonian (V.2) for $\epsilon=-1$ corresponds to the Hamiltonian constraint of a relativistic massless particle, $\mathcal{H}_{\rm{rel}}=-E^{2}+P_{i}^{2}$, where $\widetilde{\Pi}_{0}$ plays a role of “energy EE”, although this is not important for the dynamics since it is determined essentially by $H_{g\phi}\approx 0$ as discussed in Sec.I.
@@
The $(2+n)$ D-Schrödinger algebra $\mathfrak{sh}(2+n)$ consists of three sectors. The gravity sector has only gravity components $(\widetilde{\Pi}_{0},\widetilde{\Pi}_{1},\widetilde{B}^{0},\widetilde{B}^{1},\widetilde{J}_{01})$, forming together with $\widetilde{Q}_{\pm},\widetilde{Q}_{0}$ a sub-algebra $\mathfrak{sh}(2)$; the scalar-field sector is composed by $(\widetilde{\Pi}_{I},\widetilde{B}^{I},\widetilde{J}_{IJ})$, leading together with $\widetilde{Q}_{\pm},\widetilde{Q}_{0}$ to a sub-algebra $\mathfrak{sh}(n)$; and the gravity-scalar sector contains the mixed components $\widetilde{J}_{0I},\widetilde{J}_{1I}$. Therefore, in the matter decoupling limit where $\widetilde{\Pi}_{I}\to 0$, only the gravity sector remains, and the algebra reduces to the 2D-Schrödinger algebra $\mathfrak{sh}(2)$.
The remarkable point here is that this algebra $\mathfrak{sh}(2)$ is expressed in terms of $Y$-coordinates with the lapse choice (85) while the one obtained in Sec.III.2 is written in terms of $X$-coordinates under a different gauge (34). Therefore, the 2D-Schrödinger symmetry can be realized in different choices of mini-superspace coordinates and gauges. This means the covariance of the Schrödinger symmetry in the mini-superspace, which signals a robustness of the symmetry (at least, for pure-gravity cases).
@@ V.3 Generalized JNW metric and its “interior”
for $R\neq r_{-}$, where $r_{\pm}$ are constants to be determined, and $\epsilon=-1$ corresponds to $R>r_{-}$ and $\epsilon=1$ to $R<r_{-}$. We want to express the metric (81) in terms of $R$. We start with evaluating the areal radius $r$:
@@
where we have used the configuration (V.3) and (V.3) and the $\tau$-$R$ relation (V.3).
@@
where we set $\mu=1$ in the definition of $r_{0}$ (99). Although this expression is different from that of (53), we can check that $\frac{r_{0}}{2G}$ agrees with the Komar mass even in the case of $\mu\neq 1$ (see Appendix A). Note that $r_{0}$ is gauge invariant (as shown above), and so this expression of $M$ is.
@@
The spacetime volume is finite:
@@
where $\Delta t_{\infty}=\frac{l_{0}l_{s}\widetilde{\Pi}_{1}}{\bar{c}r_{0}}=\frac{l_{p}^{2}\widetilde{\Pi}_{1}}{r_{0}}$ and the definitions (99) have been used. Note that it is still non-zero after taking $\mu\to 1$, $\lim_{\mu\to 1}\mathcal{V}_{\rm 4D}=\frac{\pi l_{p}^{6}}{3l_{0}^{2}}\widetilde{\Pi}_{1}^{3}e^{-4\left(\frac{1}{\mu}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}$, but the meaning is unclear yet since the finiteness comes from the singular behavior for $\mu\to 1$ in the last factor of (V.3).
Thus, this configuration represents a Kantowski-Sachs-type closed universe for a causal-limit isotropic fluid that has initial and final singularities at $R=-r_{+},~r_{-}$ corresponding to a big bang and big crunch, respectively. Indeed, setting $n=1$ and using the scalar field $\phi_{I=1}$ as a “clock” (instead of $R$), the metric (114) agrees with the one obtained in Ref.Chiou:2008eg . The new point we find is that the closed universe and the JNW spacetime can be unified into the single expression (V.3); for the domain $r_{+}<R<r_{-}$, it represents the closed universe while for $r_{-}<R<\infty$ the JNW spacetime. In this sense, the closed universe corresponds to the “interior” of the JNW metric, although the two spacetimes are disconnected at $R=r_{-}$ where the singularity exists in both sides.
Let us conclude this section by summarizing the results. For a coupled system of gravity and $n$ massless scalar fields, we have employed the different choice of canonical variables and gauge from the ones used in the previous cases, and established the $(2+n)$-D-Schrödinger symmetry. In the vacuum case, this leads to a partial evidence for the covariance of the symmetry in the mini-superspace. Then, we have obtained the generalized JNW solution and the closed universe solution as its “interior”. They are characterized by the mass parameter $r_{0}$ and the parameter $\mu$ related to the amplitude of the scalar fields. In terms of the canonical variables, the solutions are determined by the values of the generators of the symmetry: $(\widetilde{\Pi}_{0},\widetilde{\Pi}_{1},\widetilde{B}^{0},\widetilde{B}^{1})$. Therefore, the symmetry transformation can change the values. We will discuss this point in the next section.
@@ VI Interpretation of the Symmetries under Hamiltonian Constraints
$H=\frac{1}{2}\mathcal{G}^{ab}P_{a}P_{b}+\lambda\equiv H_{0}+\lambda\approx 0,$ (117)
to reach Schrödinger symmetries. Here, $\lambda$ is a constant characterizing the dynamics, and $P_{a}$ is used to denote $\widetilde{P}_{a}$ or $\widetilde{\Pi}_{a}$ for simplicity.
@@
    Does applying a transformation $G$ to a quantity $X$ after using the constraints yield the same result as using the constraints after transforming $X$ by $G$?
@@
Let us add a little more. For (I), the physical states are restricted on a hypersurface in the phase space determined by the constraints, and some transformation $G$ may move such states outside the surface, rendering them no longer physical ones. For (II), consider, for example, transforming the ADM mass $M$ by $G=\widetilde{B}^{-}$. The definition (53) contains $\widetilde{P}_{-}$, and the mass changes by the transformation. However, if we use the constraint (43) and express $\widetilde{P}_{-}$ in $\widetilde{P}_{+}$ first, then the mass does not. For (III), physically equivalent states are represented by a single gauge orbit. If a transformation $G$ moves a gauge orbit to another one, that transformation is gauge-independent. In other words, the transformation of a gauge-invariant observable $\mathcal{O}$ by such a $G$ is gauge-invariant, where $\mathcal{O}$ commutes with the gauge constraints.
The points (I)–(III) are satisfied if a symmetry generator $G$ commutes with the gauge constraints, as shown below in an example. Actually, the transformation induced by such a generator does not alter the theory; that is, it acts as a symmetry transformation that leaves the Hamiltonian invariant.
@@ VI.1 Symmetry generating a new theory
We argue that the symmetries by non-commuting generators $G$ with the Hamiltonian constraint $H$ transform the theory.
@@
However, if $\{H_{0},G\}\neq 0$, the transformed trajectory $X^{\prime}(\tau)$ does not satisfy $H|_{P_{a}\to\alpha^{\prime}_{a}}=0$ and thus does not become a physical solution to the original theory with $H$. To interpret such a transformation as a kind of symmetry, we consider the following. For a physical solution with $(\alpha_{a},\beta^{a})$ to the original theory with $H$ (that is, $H|_{P_{a}\to\alpha_{a}}=0$), we use a generator $G$ of the Schrödinger symmetry and introduce a new characteristic constant as [^5] Here, the value of $s$ is chosen in a specific manner. See below for examples.
$\lambda_{\rm New}:=\lambda+(H_{0}-e^{s\{\cdot,G\}}\triangleright H_{0})|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}}$ (128)
@@
From a perspective of this symmetry transformation that changes $\alpha_{a}$, $\beta^{a}$, and $\lambda$ simultaneously, let us consider points (I), (II) and (III). Point (I) is achieved by adopting the idea of moving the constraint surface itself while preserving the way of transformation of a point in the phase space. For Point (II), $H=H_{0}+\lambda$ remains invariant according to this generalized transformation, and no ambiguity appears in transforming a quantity $X$.
@@
On the other hand, the transformation of gauge-invariant observables $\mathcal{O}$ can be gauge dependent. For instance, the transformation of $\mathcal{O}=J_{ab}$ by $B^{a}$ is gauge dependent, since their commutation relation is linear on $B^{a}$. Once we determine the value of $\beta^{a}$, the transformation is uniquely determined. There is room for discussion regarding the role of the gauge in this view of symmetry.
Thus, the above perspective can be considered as one interpretation of Noether’s theorem in a theory with the Hamiltonian constraint $H\approx 0$ and symmetry generators $G$ that do not commute with $H$. It is worth noting that transformations commuting with $H$ can be regarded, from this perspective, as a special case where $\lambda$ remains unchanged. Furthermore, in this interpretation, all symmetry transformations satisfying the condition (6) preserve the property (119) of leaving the time evolution operator $H_{0}$ invariant for certain choices of $N$. Therefore, this interpretation of the $\lambda$ transformation can be viewed as an extension of the symmetry perspective—where initial conditions are changed without altering the time-evolution operator $H_{0}$—to the transformation of $\lambda$.

### v2

@@ V.1 Y-coordinate in the mini-superspace
where $\eta_{ab}\equiv\mathrm{diag}(-1,1,\cdots.,1)$ is a $(2+n)\times(2+n)$ “Lorentzian” constant matrix. This contains only one coordinate $Y^{0}$ so that we can implement the canonical-transformation program easily.
@@ V.2 $(2+n)$ D-Schrödinger symmetry
$\mathcal{G}_{\phi ab}=(-\epsilon)cl_{p}\eta_{ab},~~\mathcal{G}_{\phi}^{ab}=\frac{1}{(-\epsilon)cl_{p}}\eta^{ab}.$ (93)
Here, for convenience, we multiplied 1-1 in the right hand side of (V.2) so that the Hamiltonian (V.2) for $\epsilon=-1$ corresponds to the Hamiltonian constraint of a relativistic massless particle, $\mathcal{H}_{\rm{rel}}=-E^{2}+P_{i}^{2}$, where $\widetilde{\Pi}_{0}$ plays a role of “energy EE”, although this is not important for dynamics since it is determined essentially by $H_{g\phi}\approx 0$ as discussed in Sec.I.
@@
The $(2+n)$ D-Schrödinger algebra $\mathfrak{sh}(2+n)$ consists of three sectors. The gravity sector has only gravity components $(\widetilde{\Pi}_{0},\widetilde{\Pi}_{1},\widetilde{B}^{0},\widetilde{B}^{1},\widetilde{J}_{01})$, forming together with $\widetilde{Q}_{\pm},\widetilde{Q}_{0}$ a sub-algebra $\mathfrak{sh}(2)$; the scalar-field sector is composed by $(\widetilde{\Pi}_{I},\widetilde{B}^{I},\widetilde{J}_{IJ})$, leading together with $\widetilde{Q}_{\pm},\widetilde{Q}_{0}$ to a sub-algebra $\mathfrak{sh}(n)$; and the gravity-scalar sector contains the mixed components $\widetilde{J}_{0I},\widetilde{J}_{1I}$. Therefore, in the matter decoupling limit where $\widetilde{\Pi}_{I}\to 0$ and $\widetilde{B}^{I}\to 0$, only the gravity sector remains, and the algebra reduces to the 2D-Schrödinger algebra $\mathfrak{sh}(2)$.
The remarkable point here is that this algebra $\mathfrak{sh}(2)$ is expressed in terms of $Y$-coordinates with the lapse choice (86) while the one obtained in Sec.III.2 is written in terms of $X$-coordinates under a different lapse (34). Therefore, the 2D-Schrödinger symmetry can be realized in different choices of mini-superspace coordinates and lapses. This means the covariance of the Schrödinger symmetry in the mini-superspace, which signals a robustness of the symmetry (at least, for pure-gravity cases).
@@ V.3 Generalized JNW metric and its “interior”
for $R\neq r_{-}$, where $r_{\pm}$ are constants to be determined, and $\epsilon=-1$ corresponds to $R>r_{-}$ and $\epsilon=1$ to $R<r_{-}$. We want to express the metric (82) in terms of $R$ (instead of $\tau$). We start with evaluating the areal radius $r$:
@@
where we have used the configuration (V.3) and (V.3) and the $\tau$-$R$ relation in Eq.(V.3).
@@
where we set $\mu=1$ in the definition of $r_{0}$ (100). Although this expression is different from that of (53), we can check that $\frac{r_{0}}{2G}$ agrees with the Komar mass even in the case of $\mu\neq 1$ (see Appendix A). $r_{0}$ is gauge invariant (as shown above), and so this expression of $M$ is. Note that introducing a gauge-invariant quantity $T\equiv 2\frac{G\bar{c}}{l_{s}}e^{2\left(\frac{1}{\mu}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}$, we can show that the canonical relation (57) holds for this expression of $T$ and $M$ too (in a general value of $\mu$).
@@
Furthermore, the spacetime volume is finite:
@@
where $\Delta t_{\infty}=\frac{l_{0}l_{s}\widetilde{\Pi}_{1}}{\bar{c}r_{0}}=\frac{l_{p}^{2}\widetilde{\Pi}_{1}}{r_{0}}$ and the definitions (100) have been used. Note that it is still non-zero after taking $\mu\to 1$, $\lim_{\mu\to 1}\mathcal{V}_{\rm 4D}=\frac{\pi l_{p}^{6}}{3l_{0}^{2}}\widetilde{\Pi}_{1}^{3}e^{-4\left(\frac{1}{\mu}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}$, but the meaning is unclear since the finiteness comes from the singular behavior for $\mu\to 1$ in the last factor of (V.3).
Thus, this configuration represents a Kantowski-Sachs-type closed universe for a causal-limit isotropic fluid and has the initial and final singularities at $R=-r_{+},~r_{-}$ corresponding to a big bang and big crunch, respectively. Indeed, setting $n=1$ and using the scalar field $\phi_{I=1}$ as a “clock” (instead of $R$), the metric (115) agrees with the one obtained in Ref.Chiou:2008eg . The new point we find is that the closed universe and the JNW spacetime can be unified into the single expression (V.3); for the domain $r_{+}<R<r_{-}$, it represents the closed universe while for $r_{-}<R<\infty$ the JNW spacetime. In this sense, the closed universe corresponds to the “interior” of the JNW metric, although the two spacetimes are disconnected at $R=r_{-}$ where the singularity exists in both sides.
Let us conclude this section by summarizing the results. For a coupled system of gravity and $n$ massless scalar fields, we have employed the different choice of canonical variables and lapse from the ones used in the previous cases, and established the (2+n)(2+n)D Schrödinger symmetry. In the vacuum case, this leads to a partial evidence for the covariance of the symmetry in the mini-superspace. Then, we have obtained the generalized JNW solution and the closed universe solution as its “interior”. They are characterized by the mass parameter $r_{0}$ and the parameter $\mu$ related to the amplitude of the scalar fields. In terms of the canonical variables, the solutions are determined by the values of the generators of the symmetry: $(\widetilde{\Pi}_{0},\widetilde{\Pi}_{1},\widetilde{B}^{0},\widetilde{B}^{1})$. Therefore, the symmetry transformation can change the values. We will discuss this point in the next section.
@@ VI Interpretation of the Symmetries under Hamiltonian Constraints
$H=\frac{1}{2}\mathcal{G}^{ab}P_{a}P_{b}+\lambda\equiv H_{0}+\lambda$ (118)
to reach Schrödinger symmetries. Here, $\mathcal{G}^{ab}$ is a constant invertible matrix, $\lambda$ is a constant characterizing the dynamics, and $P_{a}$ is used to denote $\widetilde{P}_{a}$ or $\widetilde{\Pi}_{a}$ for simplicity.
@@
    Does applying a transformation $G$ to a gauge-invariant observable $\mathcal{O}$ after using the constraints yield the same result as using the constraints after transforming $\mathcal{O}$ by $G$? (Here, $\mathcal{O}$ commutes with the constraints.)
@@
Let us add a little more. For (I), the physical states are restricted on a hypersurface in the phase space determined by the constraints, and some transformation $G$ may move such states outside the surface, rendering them no longer physical ones. For (II), using $H=0$, one can construct different (but weakly equivalent) expressions of $\mathcal{O}$. Applying the transformation $G$ to each generically can yield different results (even if $H=0$ is used afterward). Consider, say, transforming the ADM mass $M$ by $G=\widetilde{B}^{-}$. The definition (53) contains $\widetilde{P}_{-}$, and the mass changes by the transformation. However, if we use the constraint (43) and express $\widetilde{P}_{-}$ in $\widetilde{P}_{+}$ first, then the mass does not. For (III), physically equivalent states are represented by a single gauge orbit. If a transformation $G$ moves a gauge orbit to another one, that transformation is gauge-independent. In other words, the transformation of a gauge-invariant observable $\mathcal{O}$ by such a $G$ is gauge-invariant. See Fig.3.
Figure 3: Schematic representation of a gauge-independent transformation mapping one gauge orbit to another in the phase space. Two physically equivalent points, AA and $B$, lying on a single gauge orbit generated by $e^{s^{\prime}\{\cdot,H\}}\triangleright$, are mapped by the transformation $e^{s\{\cdot,G\}}\triangleright$. (a) The mapped points $A^{\prime}$ and $B^{\prime}$ reside on the same gauge orbit, remaining physically equivalent after the transformation. Such a transformation is gauge-independent. (b) In contrast, the mapped points $A^{\prime\prime}$ and $B^{\prime\prime}$ land on distinct gauge orbits. Here, the originally equivalent points AA and $B$ are transformed into physically distinct points. In this case, the transformation by $G$ is gauge-dependent.
The points (I)–(III) are satisfied if a symmetry generator $G$ commutes with the gauge constraints, as shown below in an example. Actually, the transformation induced by such a generator does not alter the theory; that is, it acts as a symmetry transformation that leaves the action (without fixing $N$) invariant (see Appendix B for details).
@@ VI.1 Symmetry generating a new theory
We argue that the symmetries by non-commuting generators $G$ with the Hamiltonian constraint $H$ transform the theory to a new theory.
@@
However, if $\{H_{0},G\}\neq 0$, the transformed trajectory $X^{\prime}(\tau)$ does not satisfy $H|_{P_{a}\to\alpha^{\prime}_{a}}=0$ and thus does not become a physical solution to the original theory with $H$. To interpret such a transformation as a kind of symmetry, we consider the following. For a physical solution with $(\alpha_{a},\beta^{a})$ to the original theory with $H$ (that is, $H|_{P_{a}\to\alpha_{a}}=0$), we use a generator $G$ of Schrödinger symmetry and introduce a new characteristic constant by [^6] Here, the value of $s$ is chosen in a specific manner. See below for examples.
$\lambda_{\rm New}:=\lambda+(H_{0}-e^{s\{\cdot,G\}}\triangleright H_{0})|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}},$ (129)
@@
From a perspective of this symmetry transformation that changes $\alpha_{a}$, $\beta^{a}$, and $\lambda$ simultaneously, let us consider points (I), (II) and (III). Point (I) is achieved by adopting the idea of moving the constraint surface itself while preserving the way of transformation of a point in the phase space.
For Point (II), since the generalized transformation keeps the Hamiltonian constraint as in Eq.(131), different but weakly-equivalent expressions of $\mathcal{O}$ transforms by $G$ in the same manner. For example, consider the AdS-Schwarzschild solution characterized by a set $(\alpha_{+},\alpha_{-};\beta^{+},\beta^{-})$ satisfying $H_{g}|_{\widetilde{P}_{a}\to\alpha_{a}}=0$, that is, $\frac{2}{\epsilon cl_{p}}\alpha_{+}\alpha_{-}+\lambda=0$, where $\lambda=-\frac{cl_{p}}{2l_{s}^{2}}$ from Eq.(43). By using this, the mass $M$ (53) can be expressed in various ways, say,
$M$ $=\frac{\epsilon l_{s}}{G\bar{c}^{2}}\alpha_{+}(\alpha_{+}\beta^{+}-\alpha_{-}\beta^{-})\equiv M_{1}$
$=\frac{\epsilon l_{s}}{G\bar{c}^{2}}(\alpha_{+}^{2}\beta^{+}-\frac{\epsilon cl_{p}}{2}\lambda\beta^{-})\equiv M_{2}.$
Then, apply the transformation $e^{s\{\cdot,\widetilde{B}^{-}\}}\triangleright$ (for a finite value $s$) to $M_{1}$, $M_{2}$, and $H_{g}$. According to (129), $\alpha_{-}$ and $\lambda$ transform as
$\alpha_{-}\rightarrow\alpha_{-}^{\prime}=\alpha_{-}-s,~~\lambda\rightarrow\lambda_{\rm{New}}=\lambda-\frac{2s}{\epsilon cl_{p}}\alpha_{+},$
and $\alpha_{-},\beta^{+},\beta^{-}$ do not change, which, of course, satisfies $H_{g}^{\rm New}|_{\widetilde{P}_{a}\to\alpha^{\prime}_{a}}=0$ as in Eq.(131). This leads to
$M_{1}\rightarrow M_{1}^{\prime}$ $=\frac{\epsilon l_{s}}{G\bar{c}^{2}}\alpha_{+}\left\{\alpha_{+}\beta^{+}-(\alpha_{-}-s)\beta^{-}\right\}$
$=M_{1}+\frac{\epsilon l_{s}}{G\bar{c}^{2}}\alpha_{+}\beta^{-}s,$
$M_{2}\rightarrow M_{2}^{\prime}$ $=\frac{\epsilon l_{s}}{G\bar{c}^{2}}\left\{\alpha_{+}^{2}\beta^{+}-\frac{\epsilon cl_{p}}{2}\left(\lambda-\frac{2s}{\epsilon cl_{p}}\alpha_{+}\right)\beta^{-}\right\}$
$=M_{2}+\frac{\epsilon l_{s}}{G\bar{c}^{2}}\alpha_{+}\beta^{-}s.$
Therefore, $M_{1}^{\prime}=M_{2}^{\prime}$ holds, and the way of the transformation does not depend on the expression of $M$. Note that if $\lambda$ were not transformed simultaneously, we would have $M^{\prime}_{1}\neq M^{\prime}_{2}$.
@@
On the other hand, the transformation of gauge-invariant observables $\mathcal{O}$ can be gauge dependent. For instance, the transformation of $\mathcal{O}=J_{ab}$ by $B^{a}$ is gauge dependent, since their commutation relation is linear on $B^{a}$. (This corresponds to the case (b) in Fig.3.) Once we determine the value of $\beta^{a}$, the transformation is uniquely determined. There is room for discussion regarding the role of the gauge in this view of symmetry.
Thus, the above perspective can be considered as one interpretation of Noether’s theorem in a theory with the Hamiltonian constraint $H\approx 0$ and symmetry generators $G$ that do not commute with $H$. It is worth noting that transformations commuting with $H$ can be regarded, from this perspective, as a special case where $\lambda$ remains unchanged. Furthermore, in this interpretation, all symmetry transformations satisfying the condition (6) preserve the property (120) of leaving the time evolution operator $H_{0}$ invariant for certain choices of NN. [^9] The invariance of $H_{0}$ for a specific choice of $N$ corresponds to leaving the gauge-fixed action (165) invariant. See Appendix B. Therefore, this interpretation of the $\lambda$ transformation can be viewed as a generalization of the standard dynamical symmetry.